(5) (20p) Find and classify the critical points of f(x, y) = 7x - 8y + 2xy - x + y®
Answer #1

f(a,y) = 78-8y + 2xy - arty? fx=7-0724-2840 = 7+2y-22; fy=+8+2a+zy. 7-8 +22+2y fxx - I (fx) = -2; fyy = 2 (fy)=2 27=24+ To fi